OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-2, 4) and is parallel to x - 2y = 6

OpenStudy (anonymous):

-2y=-x+6 y=(1/2)x-3 4=(1/2)(-2)-b 4=-1 - b 5=b y=(1/2)x+5 Slope intercept form -(1/2)x+y=5 standard form

OpenStudy (anonymous):

-(1/2) cant be an answer..

OpenStudy (anonymous):

but that 1/2 is the slope

OpenStudy (anonymous):

So what is just the standard form?